OpenStudy (blackops2_143):

which ratio is equivalent to 18/24 a. 5/8 b. 2/3 c. 10/12 d. 3/4

OpenStudy (blackops2_143):

is it b?

OpenStudy (skullpatrol):

Nope.

OpenStudy (skullpatrol):

18 = 6 X 3 24 = 6 X 4

OpenStudy (skullpatrol):

$$\Huge \dfrac{18}{24}=\dfrac{6 \times 3}{6 \times 4}$$

OpenStudy (blackops2_143):

thank you

OpenStudy (skullpatrol):

Thanks for learning :-)